Find The Sum or Difference.
25.-6 2/3 - 4 3/5
Any help with this equation would be GREATLY appreciated.

Answer by stanbon(75887)

-6 2/3 - 4 3/5
= -10 -(2/3+3/5)
= -10 -(10+9)/15
= -10 - 19/15
= -11 4/15
=============
Cheers,
Stan H.